OpenStudy (anonymous):

PLEASE help I have been stuck on this for hours! If a person starts to learn y items in x hours as given by y=21(x^2)^1/3 find the rate of learning for a person at the end of three hours. I start out like this 1/3(21)(x^2)^-3/2 so 7(x^2)^-3/2 but now I am stuck and I don't know if what I did above was correct. Please help!

OpenStudy (aaronq):

take the derivative which is the rate or slope (whatever you wanna call it)

OpenStudy (aaronq):

then plug in x=3 since 3 hours

OpenStudy (anonymous):

Yes, but how do you solve the derivative?

OpenStudy (aaronq):

www.wolframalpha.com they have step by step

OpenStudy (anonymous):

Hint: You need to simplify first. y=21(x^2)^1/3=21x^2/3 unless i am reading this wrong. Then this becomes a very simple derivative. If you cant solve it let me know.

OpenStudy (anonymous):

Here is a picture of the equation if that would help

OpenStudy (anonymous):

The derivative of 21X^2/3=2/3(21)x^-1/3=14X^-1/3 0r 14/x^1/3 so at x=3, this is goint to equal 14/3^1/3